Grade 12Algebra

Let n be the number of points having rational coordinates at a constant distance from the point (0,3^1/2), then
a.n>2
b.n=1
c.n?=2
d.n=1

anjani
10 Years agoGrade 12

2 Answers

Vikas TU
10 Years ago
let (x,y) be point at constant distance from (0,root3).
then, acc. to question,
distancee from point (0,root3) to (x,y) should be equal to root3.
.i.e
root(x^2 + ((y-root3)^2) = root3
Or
x^2 + (y-root3)^2  =  3
x^2 + y^2 + 3 – 2yroot3 = 3
x^2 + y^2 – 2yroot3 = 0
Now comparing rational and irrational terms for distinct coordinates,
.i.e.
x^2 + y^2 = 0     and     yroot3 = 0 = > y = 0
at y = 0 
x = 0.
hence (0,0) is the only point .
hence n = 1
 
HARISH SAHU
9 Years ago

Let A(x1, y1) and B (x2, y2) be two points at equal distance d from (0, 31/2)

then 

(x1-0)2 + (y1-0)2 = d2..............(1)

(x2-0)2 + (y2-0)2 = d2..........(2)

from equation (1) and (2)

x12 - x22 +(y1 - 31/2) - (y- 31/2) =0

x12 - x22 + (y12 - y22) -2*31/2 (y-  y2) =0

comparing rational and irrational parts on both sides

y1 - y2 =0   => y1 = y2 = a (suppose)

x12 -x22 + (y12 - y22) =0 

x12 = x22 =b2

x= x\pm b

so points are (\pm b,a)

so n\leq 2